# A new criterion for interpreting acoustic emission damage signals in condition monitoring based on the distribution of cluster onsets

### v1

@@ 2.2 Case 2: Without a priori on onsets
We can simplify Eq. 7 by dividing by $\log_{2}K$ leading to
@@
where $E_{\textrm{norm}}\in[0,1]$ is the normalized entropy of ${\mathbf{p}}_{\textrm{est}}$. Therefore, for a given $K$, minimizing the initial criterion in Eq. 3 divided by $\log_{2}K$ is equivalent to maximizing the entropy of onsets. A possible criterion to sort subsets is thus:
@@
Example 2.
[…]
Remark 1.
One important point to observe is that the proposed CVI only requires the partition, not the features conversely to shape-based CVI. Therefore, according to [25], the proposed criterion belongs to the family of “external” CVI whereas shape-based are said “internal”. However, the proposed criterion is a particular case because external CVI actually require a complete reference partition to be compared with, meaning that the ground truth must be known for all feature vectors, whereas the proposed CVI only requires the onsets of clusters.
@@ 2.3 Proposed algorithm
The algorithm based on onsets for evaluating clustering results of AE data is presented in Alg. 2 (Appendix B). In this algorithm, $Q_{\textrm{proposed}}^{\textrm{ONSETS}}(S,K)$ used in line 10 is given by Eq. 9 and has to be maximized. In line 12, partitions are sorted according to the value of this criterion and a few dozens of the best ones are selected. For example, the value of the first percentile of the values of $Q_{\textrm{proposed}}^{\textrm{ONSETS}}$ can be used to determine the number of partitions to be selected. For each selected partition, the onsets are determined (line 15) by simply finding the time of the first occurrence of each cluster. These steps are done for each number of clusters, separately.
Finally, onsets are stacked independently on the number of clusters and the result is presented by means of a histogram that approximates the distribution of onsets. Results of the accumulation of onsets are presented to the end-user who visually determines the position of onsets by analyzing the peaks at some places in the histogram. As in interactive clustering [26], based on the histogram, the end-user can then either reject, accept, or be ignorant of the result, triggering a re-clustering of the data, for example using different features or another clustering method.
The histogram can be interpreted as follows:
  • High peaks mean that varying the number of clusters and features lead to similar locations of onsets. According to the application, this can reflect a robustness of the clustering method or its inability to decompose the data set into clusters.
  • Many small peaks mean that a change of the number of clusters implies a systematic change of the location of onsets. In that case, the clustering method is not robust or the features are not relevant.
  • In real AE data sets, it is expected to get high peaks located at various locations surrounded by smaller peaks. It means that some clusters are found with different onsets but varying the number of clusters or the features implies a variability on the onsets.
  • Since the first cluster starts at 0 by construction (Eq. 1) for all features and all clusters, the first peak will be the highest and represents the number of combinations selected to plot the histogram.
The location of the peaks can be helpful for the end-user to conclude that some damages potentially start or grow from these locations or if the clustering is meaningful or not.
@@ 3 Experiments
Three applications have been considered. Table 1 present the details about their parametrisation such as the number of features per subset, the possible number of clusters and the clustering method. For each application, the results of the proposed onset-based CVI (Alg. 2) and CVI based on clusters’ shape [10] (Algorithm 1) are presented.
@@
Bolted metallic 1 19 4 $\{2,3,4\dots 10\}$ Yes (7 classes) Kmeans
@@ 3.1.1 Setup and data description
The benchmark data set ORION-AE is presented in [27]. In this data set, the true labels are provided which allowed us to evaluate the performance of the proposed criterion and compare it with the standard solution. The data were collected on a setup designed to reproduce the loosening phenomenon observed in aeronautics, automotive or civil engineering structures where parts are assembled together by means of bolted joints (Figure 4).
@@
The features are freely available at the following link [27]: https://drive.google.com/drive/folders/1H413RxYu4ya7YMEgF_lTh_fHr7flvvOO?usp=sharing. They are common features used in the AE community [28]: Rise time, counts, counts to peak, duration, MARSE and absolute energies, amplitude, average signal level, signal strength, RMS, reverberation frequency, initiation frequency, average frequency, partial power in [0,200], ]200,400], ]400,600], ]600,1000] kHz, peak frequency and centroid frequency. The set of feature vectors was post-filtered using a 31-sample moving median applied to each dimension in order to ensure temporal coherence, and one point over three were kept in order to reduce the amount of points (there are about 1200 cycles per tightening levels in the original data set).
3.1.2 Parametrization of algorithms
Algorithms 1 and 2 were used on every combination of $4$ features within the set of $19$ mentioned above, so leading to $3876$ combinations for each $K=3,4\dots 10$ value with a total of $31008$ combinations. Kmeans clustering method was run 5 times for every combination and the result leading to the minimum sum of squared euclidean distances over all points and clusters was selected. If an empty cluster was obtained, the corresponding subset of features was not used. For Algorithm 1 (based on shape), the voting scheme proposed in [10] was applied. We used three internal CVI implemented in Matlab’s Statistical and Machine Learning Toolbox, namely: Silhouette, Calinski-Harabasz and Davies-Bouldin. We used the same 25 points voting scheme as in [10]. Algorithm 1 was applied for each $K$ independently in order to sort the subsets of features and select the best one. For Algorithm 2 (onset-based), the $20$ best partitions were selected for each number of clusters using the onset-based CVI.
@@ 3.1.3 Link between clusters onsets and change in tightening levels
We evaluated whether the clusters’ onsets coincide with the changes of tightening levels. For that, we represented the distribution of the estimated onsets using histograms. This was done in all partitions found by the previous algorithms and for all numbers of clusters ($K=3\dots 10$). Histograms depict the ground truth with blue stairs and the vertical dotted lines represent the instant of the true changes in the tightening level. This ground truth is available in [27]. The stairs show the duration of each level, equal to $10.0,10.0,10.0,9.0170,10.0,10.0$ and $10.0$ seconds for, respectively, $60,50,40,30,20,10$ and $5$ cNm.
Histograms of onsets for shape-based and onset-based CVI are presented in Figure 6 with a bin size of $0.5$ s. These results only differ by the criterion since the same partitions obtained by the Kmeans method were used to evaluate both criteria in the same conditions. With the shape-based approach (Fig. 7(a)), we can observe three missing onsets (no vertical bar at 10, 30 and 60 s corresponding to 50, 30 and 5 cNm), one delayed onset (20 s / 40 cNm) and one clear false positive (at around 44 s). Based on this histogram, an end-user is likely to select 4 clusters, with one wrong and 3 missed tightening levels.
(a) Shape-based CVI with the best vote
(b) Proposed onset-based CVI with 20 best partitions
Figure 6: Validation of the clustering provided by a Kmeans using the voting scheme and onset-based criteria.
Figure 7(b) shows that all changes in tightening levels, except the last one, are detected by the onset-based CVI. Four primary peaks distinctly appear, located at expected times corresponding to changes in tightening levels at 60 cNm (0 s), 50 cNm (10 s), 40 cNm (20 s), 20 cNm (40 s), 10 cNm (50 s). The same partitions were used in both cases which means that only the CVI makes the difference. The peak at 30 s is well distinguishable with 17 occurrences. The secondary peaks (lower occurrence) are globally distributed around the main ones and represent the variability on onsets in the 20 partitions. Onset at 60 s corresponding to 5 cNm was not found because of a mixing made by the clustering with the data in the previous tightening level (50 s / 10 cNm).
@@ 3.1.4 Link between clusters onsets and clusters overall quality
We evaluated whether the content of clusters (over all tightening levels) coincide with the ground truth. For that, the 20 best partitions provided by each criterion were evaluated using the Rand Index (RI) [29]. RI is a value in $[0,1]$ reflecting the quality of clustering, where “$1$” means that the assignment of clusters made by clustering is correct for all feature vectors. The number of points in each partition is 4684.
Results are presented in Figure 7 using boxplots. The red line is the median value, and the box is defined by the minimum and maximum whisker length (W=1.5 as the default in Matlab) so that RI values are considered as outliers if they are larger than Q3+W*(Q3-Q1) or smaller than Q1-W*(Q3-Q1), where Q1 and Q3 are the 25th and 75th percentiles, respectively. The outliers are represented with red crosses. The figure shows that, overall, the median RI values are globally close. If we consider values around the ground truth ($K_{true}=7$), the main difference is the width of the boxes which is largest for the shape-based CVI. This is confirmed in Table 2 which is a summary of the RI values for $K=6,7,8$. The table shows that the minimum RI values are higher for the onset-based with an improvement of $4\%$ to $7\%$, while the maximum values are similar. We can also see that $4$ values are found as outliers below the minimum (and 11 for all values of $K$) for the shape-based criterion, representing $10\%$ of the results. Conversely, no outlier was detected with the onset-based criterion for $K=6,7,8$ (and $3$ for all values of $K$), demonstrating its higher robustness to the change of features.
(a) Partitions selected by shape-based CVI
(b) Partitions selected by onset-based CVI
Figure 7: Performance evaluated by the Rand Index
Nb cluster/perf min max median nb outliers
@@ 3.1.5 Case of a data set with non-uniformly distributed onsets
In ORION-AE data set, the tightening levels have approximately the same duration which is a special case. In this section, we propose to evaluate the proposed criterion when the duration in each tightening level is not similar. For that, we selected only a part of the data in each level so that the duration of the levels follows a non linear evolution. Practically, it means that the metallic plates remain tightened for a longer period for a given level compared to the subsequent levels. For $60$, $50$, $40$, $30$, $20$, $10$ and $5$ cNm, the first seconds of data were kept as follows: $10\,s$, $8.5\,s$, $7\,s$, $5\,s$, $3\,s$, $2\,s$ and $1.5\,s$, and the remaining data were discarded.
The results are reported in Figure 8. We can observe that the proposed onset-based CVI outperforms the standard approach. As in the previous case, only the last onset is not discovered by the onset-based CVI. Conversely, the shape-based criterion found three onsets. For 50 cNm (10 s) and 40 cNm (20 s) only one occurrence is found which not enough to distinguish these onsets from the false positives ones around.
(a) Shape-based CVI with the best vote
(b) Proposed onset-based CVI with 20 best partitions
Figure 8: Case of non-uniform onsets: Validation of the clustering provided by a Kmeans using the voting scheme and onset-based criteria.
@@ 3.1.6 Summary
When averaged over 20 partitions, each made of more than 4800 points, we observed that the quality of clusters found by the onset-based criteria is improved by $4\%$ to $7\%$. An improvement can also be observed with the proposed onset-based CVI in terms of accuracy and robustness to changes in features and number of clusters. In terms of onsets, a distinct improvement was observed with the onset-based CVI which makes this criterion suitable for structural health monitoring purposes. The results illustrate the interest of this kind of criterion to detect changes in AE data by clustering.
3.2 Data set #$\#2$: Monitoring of an electrical discharge machine during microdrilling
This application is about monitoring of a manufacturing process called Micro Electrical Discharge Machining (Micro-EDM) used for drilling micro-diameter and deep holes. The goal of AE clustering is here to get insights about the process, and to evaluate if it is possible to detect when the electrode has reached the end of drilling and if the process proceeds similarly between two drillings.
@@ 3.2.1 Materials and method
Micro-EDM relies on electrical discharges emitted by a tool-electrode and used to remove material in a work piece hosed by a dielectric (Fig. 8(a)). An electric arc is formed between the conducting part and the electrode which are not in contact, making it possible to machine very hard conductive materials with a very high flexibility in terms of shapes machined [30]. Both the feed rate of the tool-electrode and the sequence of discharges, necessary to remove pieces of material to reach a reference in depth, are controlled by a proprietary software. Moreover, during the process, both the work piece and the tool-electrode get damaged and it is difficult and time-consuming to evaluate precisely and in real-time their degree of wear. We therefore used the acoustic emission technique to collect data during drilling with the aim to get insights about the process. More specificially, the goal is to evaluate whether the chronology of acoustic emission signals helps in monitoring the progression of the electrode until the reference depth.
Experiments were conducted on a SARIX EDM machine. The particularity of this machine is the auto-regulation of input parameters for the sparkles generation. A $0.3$ mm diameter tubular electrode was used to drill a $2$ mm hole opening into a $0.3$ mm thick cavity. The drilled material is a Ni-based superalloy. Figure 8(b) shows the set-up on the EDM machine. The camera is synchronized with the start of the process to record the moment when the electrode starts drilling. The camera is then used to visually evaluate when a sparkle appears in the cavity corresponding to the opening of the hole ($3^{\textrm{rd}}$ image in Figure 8(a)). This event is noted on the figures below since we expect a particular behavior on the AE signals and clusters.
AE streaming were acquired at a sampling frequency of 5 MHz, using a "micro-80" sensor made by Mistras-EuroPhysical Acoustics, with an operating frequency range in 200-900 kHz. Acquisition of stream data was made using a Picoscope from Picotech. The AE hits were detected using the same method as in application #1 (ORION-AE) and with similar features except the partial powers, considered in the following intervals: in 0-20 kHz, 20-100 kHz, 100-200 kHz, 200-300 kHz, 300-400 kHz, 400-500 kHz, 500-600 kHz, 600-800 kHz, 800-1000 kHz. A total number of 24 features were considered.
(a) Principle of EDM drilling.
(b) Illustration of the AE acquisition set-up.
Figure 9: EDM principle and instrumentation [30].

### v3

@@ 2.2 Case 2: Without a prior on onsets
We can simplify Eq. 7 by dividing by $\log_{2}K$, leading to
@@
where $E_{\textrm{norm}}\in[0,1]$ is the normalised entropy of ${\mathbf{p}}_{\textrm{est}}$. Therefore, for a given $K$, minimising the initial criterion in Eq. 3 divided by $\log_{2}K$ is equivalent to maximising the entropy of onsets. A possible criterion to sort subsets is thus:
@@
Example 2.2.
Let us continue the previous example, with $T=100$ and $K=3$ we can define three onsets distributed uniformly in order to define ${\mathbf{p}}_{\textrm{true}}$, located at $t_{1}^{\textrm{true}}=0$, $t_{2}^{\textrm{true}}=T/3=33.33$, $t_{3}^{\textrm{true}}=2T/3=66.66$, with all $\Delta_{k}=33.33/T$. Therefore ${\mathbf{p}}_{\textrm{true}}={\mathbf{p}}_{\textrm{uniform}}=[0.33,0.33,0.33]$ (at two digits of precision, the sum must be one). Using ${\mathbf{p}}_{\textrm{est}}$ of the previous example, we get $KL=0.0454714$ (Eq. 3) with $CE=1.585$ and $E=1.539$. The maximum of $E$ for $3$ clusters is $\log_{2}(3)=1.585$ therefore $E_{\textrm{norm}}=1.539/1.585=0.971$.
Remark 2.1.
One important point to observe is that the proposed CVI only requires the partition, not the features, conversely to a shape-based CVI. Therefore, according to [57], the proposed criterion belongs to the family of “external” CVIs whereas shape-based CVIs are referred to as “internal”. However, the proposed criterion is a special case because external CVIs actually require a complete reference partition to be compared with, meaning that the ground truth must be known for all feature vectors, whereas the proposed CVI only requires the onsets of clusters.
Remark 2.2.
Our criterion can be used in an online setting for SHM applications, provided $T$ is known. $T$ represents, for example, the maximum number of cycles for fatigue testing, or the maximum crack size. Since in an online setting, $K$ is assumed to evolve, we need to iteratively store the onsets. The list of onsets will be updated with a new value (appending a new element in a vector) when a new cluster is created. Then, Eq. 1 and 2 can be computed, followed by Eq. 9 (for example, without prior).
@@ 2.3 Proposed algorithm
The criterion presented in the previous section is implemented within a broader procedure that incorporates the steps of the feature-based clustering method outlined in Figure 1. This procedure proceeds as follows.
First, a standard clustering method (in this case, K-means) is applied to the dataset, assuming a range for the number of clusters and possible feature subsets. Unlike other methods, for each partition an additional characteristic, called the cluster onset, is defined as the first occurrence of that cluster in the dataset. The criterion, given by Eq. 5 or Eq. 9, depending on whether prior information on the onsets is available, is then calculated for each partition. The best partitions are selected based on the maximisation of this criterion (e.g., the top percentile of the values of $Q_{\textrm{proposed}}^{\textrm{ONSETS}}(S,K)$).
Finally, the onsets of each cluster in the selected partitions are extracted and compiled into a histogram representing the cumulative distribution of onsets across all selected partitions. Similar to interactive clustering [4], the end-user can analyse this histogram, and based on the positions and amplitudes of peaks, they can choose to either accept, reject, or further refine the results—potentially triggering a re-clustering process by using different features or alternative clustering methods.
An algorithm detailing this procedure is provided in Alg. 1.
Inputs:
    Matrix of features $\mathbf{Y}$ with size $N\times d$
    Range of number of clusters $[K_{\textrm{min}},K_{\textrm{max}}]$
    Subsets of features $\mathcal{S}$
Outputs:
    Criterion
    Histogram of onsets
1
2for $K\in[K_{\textrm{min}},K_{\textrm{max}}]$ do
3       for $S\in\mathcal{S}$ do
5             $[P(S,K),\theta(S,K)]\leftarrow$ Clustering($\mathbf{Y}(:,S)$, $K$);
6
7             // Compute quality according to how onsets are distributed;
9
10       end for
11
15
16       // Find onsets of the selected partitions;
17       $\textrm{Onsets}(K)=\textrm{Find\_onsets}(F(K))$;
18
19 end for
24
25
The interpretation of the histograms can vary across applications and specificity of the dataset. The location of the peaks can be helpful for the end-user to conclude that some damages potentially start or grow from these locations. It can also lead to the conclusion that the extracted features or the clustering method are meaningful or not. Examples are provided in the following sections.
@@ 3 Experiments
Three applications have been considered. Table 1 presents the details of their parameterization, including the number of features per subset, the possible number of clusters, and the clustering method. For each application, the results of the proposed onset-based CVI (Alg.1) and the shape-based CVI [53] (Alg. 2) are presented.
@@
Bolted metallic 1 19 4 $\{2,3,4\dots 10\}$ Yes (7 classes) K-means
@@ 3.1 Dataset #$\#1$: Monitoring of a bolted metallic structure
In this section, three state-of-the-art shape-based CVIs used in the AE community, namely the Silhouette, Davies-Bouldin, and Calinski-Harabasz CVIs, are used to sort the partitions with the voting scheme [53] and compared to the proposed onset-based CVI to provide a benchmark. It is important to note that these CVIs are adapted for hard clustering methods like K-means. Soft clustering techniques such as Gaussian mixture models (GMM) or fuzzy C-means (FCM), which assign probabilistic memberships, may not align well with these metrics. Therefore, K-means is used in both cases as well as because it is a widely adopted method in AE literature.
@@ 3.1.1 Setup and data description
The benchmark dataset ORION-AE is presented in [48]. In this dataset, the true labels are provided, which allowed us to evaluate the performance of the proposed criterion and compare it with the standard solution. The data were collected on a setup designed to reproduce the loosening phenomenon observed in aeronautics, automotive, or civil engineering structures where parts are assembled together by means of bolted joints (Figure 4).
@@
The features are freely available at the following link [48]: https://drive.google.com/drive/folders/1H413RxYu4ya7YMEgF_lTh_fHr7flvvOO?usp=sharing. They are common features used in the AE community [27]: Rise time, counts, counts to peak, duration, MARSE and absolute energies, amplitude, average signal level, signal strength, RMS, reverberation frequency, initiation frequency, average frequency, partial power in [0,200], [200,400], [400,600], [600,1000] kHz, peak frequency, and frequency centroid. The set of feature vectors was post-filtered using a 31-sample moving median applied to each dimension in order to ensure temporal coherence, and one point in three were kept in order to reduce the amount of points (there are about 1200 cycles per tightening level in the original dataset).
3.1.2 Parametrisation of algorithms
Algorithms 1 and 2 were used on every combination of $4$ features within the set of $19$ mentioned above, resulting in $3876$ combinations for each $K=3,4\dots 10$ value with a total of $31008$ combinations. K-means clustering was run 5 times for every combination and the result leading to the minimum sum of squared Euclidean distances over all points and clusters was selected. If an empty cluster was obtained, the corresponding subset of features was not used. The set of partitions obtained was used in both criteria. For Algorithm 2 (based on shape), the voting scheme proposed in [53] was applied with three internal CVIs implemented in Matlab’s Statistical and Machine Learning Toolbox: Silhouette, Calinski-Harabasz, and Davies-Bouldin. Algorithm 2 was applied for each $K$ independently in order to sort the subsets of features and select the best one. For Algorithm 1 (onset-based), the $20$ best partitions were selected for each number of clusters using the onset-based CVI.
@@ 3.1.3 Link between clusters onsets and change in tightening levels
We evaluated whether the clusters’ onsets coincide with the changes of tightening levels. For that, we represented the distribution of the estimated onsets using histograms. This was done in all partitions found by the previous algorithms and for all numbers of clusters ($K=3\dots 10$). Histograms depict the ground truth with blue stairs and the vertical dotted lines represent the instant of the true changes in the tightening level. This ground truth is available in [48]. The stairs show the duration of each level, equal to $10.0,10.0,10.0,9.0170,10.0,10.0$, and $10.0$ seconds for $60,50,40,30,20,10$, and $5$ cNm, respectively.
Histograms of onsets for shape-based and onset-based CVIs are presented in Figure 6 with a bin size of $0.5$ s. These results only differ by the criterion since the same partitions obtained by the K-means method were used to evaluate both criteria in the same conditions. With the shape-based CVI (Fig. LABEL:figB142a, we can observe three missing onsets (no vertical bar at 10, 30, and 60 s corresponding to 50, 30, and 5 cNm), one delayed onset (20 s / 40 cNm) and one clear false positive (at around 44 s). Based on this histogram, an end-user is likely to select 4 clusters, with one wrong and 3 missed tightening levels.
Figure 6: Validation of the clustering provided by a K-means using the voting scheme and onset-based criteria.
In comparison, Figure LABEL:figB142b shows that all changes in tightening levels, except the last one, are detected by the onset-based CVI. Four primary peaks distinctly appear, located at expected times corresponding to changes in tightening levels at 60 cNm (0 s), 50 cNm (10 s), 40 cNm (20 s), 20 cNm (40 s), 10 cNm (50 s). The same partitions were used in both cases which means that only the CVI makes the difference. The peak at 30 s is clearly distinguishable with 17 occurrences. The secondary peaks (lower occurrence) are globally distributed around the primary peaks and represent the variability on onsets in the 20 partitions.
High peaks mean that changing the number of clusters and features does not modify the location of these onsets. However, as demonstrated by Sawan et al. [54], we expect that increasing the number of clusters gives different insights, which is clearly not the case with the shape-based CVI. Moreover, we observe that the number of small peaks for the onset-based CVI is larger, explained by the fact that 20 times more partitions were used, adding variability to the results. We observe that the average level of these peaks is below $10$, representing less than $8\%$ of values of highest peaks. This variability highlights the difficulty in detecting the tightening level at 40 cNm, which may be represented by more than one cluster. The high number of small peaks at the highest tightening levels (left-hand side) can be attributed to the challenge in accurately quantifying bolts loosening from vibration data [38, 17, 13].
@@ 3.1.4 Link between clusters onsets and clusters overall quality
We have shown above that the onset-based CVI provides more relevant onsets. To further evaluate its performance, we assess whether the clusters generated coincide with the ground truth. For this, the 20 best partitions provided by each criterion were evaluated using the Rand index (RI) [49]. The RI is a value in the range $[0,1]$, where "$1$" indicates a perfect match between the clustering assignments and the ground truth. The number of points in each partition is 4684, based on which the RI is computed. The results are presented in Figure 7 using boxplots whereby the red line represents the median value, the boxes indicate variability based on the 25th and 75th percentiles, and the outliers are marked by red crosses.
We observe that the median RI values are generally close, indicating that, on average, both criteria provide similar cluster assignments across 4684 points. Note that the onsets are obscured in these results: only the histograms shown earlier allow for an assessment of their precise positions.
Concerning the variability across partitions, it differs between the two criteria. When considering values corresponding to the ground truth ($K_{true}=7$), the onset-based CVI demonstrates better consistency, with less variability in results compared to the shape-based CVI. This observation is confirmed in Table 2, which summarizes the RI values for $K=6,7,8$. The minimum RI values show a 4% to 7% improvement for the onset-based CVI, while maximum values are comparable for both criteria. This suggests that the worst partitions identified by the onset-based criterion are more relevant than those produced by the shape-based CVI. Additionally, four outlier values below the minimum were found for the shape-based CVI (11 across all $K$ values), representing 10% of the results. In contrast, no outliers were detected for the onset-based criterion at $K=6,7,8$, and only three for all $K$ values. These results demonstrate its superior robustness in response to feature changes.
Figure 7: Performance comparison evaluated by the Rand index.
# cluster/perf min max median # outliers
@@ 3.1.5 Case of a dataset with non-uniformly distributed onsets
In the ORION-AE dataset, the tightening levels have approximately the same duration, which is a special case. In this section, we propose to evaluate the proposed criterion when the duration in each tightening level differs. The dataset becomes imbalanced. For that, we selected only a part of the data in each level so that the duration of the levels follows a non linear evolution. Practically, it means that the metallic plates remain tightened for a longer period for a given level compared to the subsequent levels. For $60$, $50$, $40$, $30$, $20$, $10$ and $5$ cNm, the first seconds of data were kept as follows: $10\,s$, $8.5\,s$, $7\,s$, $5\,s$, $3\,s$, $2\,s$ and $1.5\,s$, and the remaining data were discarded.
The results are reported in Figure 8. We can observe that the proposed onset-based CVI outperforms the standard approach by selecting relevant partitions for which onsets are located as expected. As in the previous case, only the last onset is not discovered by the onset-based CVI. Conversely, the shape-based CVI only found three clear onsets. For 50 cNm (10 s) and 40 cNm (20 s) only one occurrence is found, which is not enough to distinguish these onsets from the false positives.
The difference in behaviour of both criteria between this test (irregular timing of damage mechanisms) and the previous one (regular timing of damage mechanisms) is the size of the clusters. Since we shortened the duration of each tightening level, there are fewer points in the classes. Consequently, the distribution of the features within the classes has changed, making the clustering task more difficult. Some classes becomes smaller relatively to others due to this change of distribution which affects both the intra-cluster compactness and inter-cluster separability, therefore degrading the values of the shape-based CVIs.
Figure 8: Case of non-uniform onsets: Validation of the clustering provided by a K-means using the voting scheme and onset-based criteria.
@@ 3.1.6 Summary
When averaged over 20 partitions of ~4600 points, we observed that the quality of clusters found by the onset-based criteria is improved by $4\%$ to $7\%$, while using the same partitions as inputs, i.e. all else being equal. With histograms, a distinct improvement was observed on the positioning of onsets with the proposed CVI, which makes this criterion suitable for structural health monitoring.
3.2 Dataset #$\#2$: Micro-drilling by electrical discharge machining
This application is about monitoring of a manufacturing process known as micro-electrical discharge machining (micro-EDM) used for drilling micro-diameter and deep holes. The goal of AE clustering is to gain insights about the process, and to evaluate if it is possible to detect when the electrode has reached the end of drilling and if the process proceeds similarly between two drilling events.
@@ 3.2.1 Materials and method
Micro-EDM relies on electrical discharges emitted by a tool-electrode to remove material from a work-piece immersed in a dielectric fluid (Fig. LABEL:edm1). An electric arc forms between the conducting part and the electrode, which are not in physical contact, enabling the machining of very hard conductive materials with a high degree of flexibility in terms of shape. Both the feed rate of the tool-electrode and the sequence of discharges, which are necessary to remove material and achieve a specified depth, are controlled by proprietary software. During this process, both the work-piece and the tool-electrode experience wear, which is challenging and time-consuming to evaluate precisely and in real-time. Therefore, we used AE to collect data during drilling to gain insights into the process. Specifically, the goal is to assess whether the chronology of AE signals can aid in monitoring the progression of the electrode to the reference depth.
Experiments were conducted on a SARIX EDM machine, which features auto-regulation of input parameters for spark generation. A $0.3$ mm diameter tubular electrode was used to drill a $2$ mm hole into a $0.3$ mm thick cavity. The material being drilled is a Ni-based superalloy. Figure LABEL:edm2 illustrates the setup on the EDM machine. The camera, synchronised with the start of the process, records the moment when the electrode begins drilling. It is then used to visually assess when a spark appears in the cavity, corresponding to the hole opening (third image in Figure LABEL:edm1). This event is marked in the figures below as we anticipate specific behaviour in the AE signals and clusters. AE data streams were acquired at a sampling frequency of 5 MHz using a "micro-80" sensor made by Mistras-EuroPhysical Acoustics, with an operating frequency range of 200-900 kHz. Data acquisition was performed using a Picoscope from Picotech. The AE hits were detected using the same method as in application #1 (ORION-AE) and similar features, except for the partial powers, which were considered in the following intervals: [0 20], [20 100], [100 200], [200 300], [300 400], [400 500], [500 600], [600 800], and [800 1000] kHz, resulting in a total of 24 features.
Figure 9: EDM principle (left) and instrumentation (right).